# Stability theory over toroidal or Novikov-type base and canonical modifications

### v1

@@ 3.2. Higher $\Theta$-stable reduction theorem
We take a regular simplicial subcone $\sum_{i=1}^{r}\mathbb{R}_{\geq 0}v^{\prime}_{i}$ of $\tau^{\prime}$, where $\{v^{\prime}_{i}\}_{i}\cup(\vec{0},1)$ is a basis of of $N$, which still contains $\xi^{\prime}$. We also take the cone which contains $(\xi,[c^{\prime}-\epsilon,c^{\prime}])$ for $0<\epsilon\ll 1$ and denote it as $\widetilde{\tau}^{\prime}$. Write $v^{\prime}_{i}=(v_{i},\frac{1}{m_{i}})$ with $m_{i}\in\mathbb{Z}_{>0}$ and set $m:={\rm lcm}\{m_{i}\}$.
Take an extension of discrete valuation ring $R\to R^{\prime}$ with ramification degree $m$, which is obtained as a localization of $R[x]/(x^{m}-\pi)$ at $(x)$. Note that $\{\frac{m_{i}v^{\prime}_{i}}{m}\}_{i}$s together with $(\vec{0},\frac{1}{m})$ gives a basis of a regular cone of $N\oplus\frac{1}{m}\mathbb{Z}$. From the construction, (base change of) $\tilde{f}$ is extendable to $\widetilde{\rm ST}_{R^{\prime}}(\tau^{\prime})\to M$ and ${\overline{\rm ST}}_{R^{\prime}}(\tau^{\prime})\to\mathcal{M}$ and the desired properties are satisfied from the construction and finish the proof of (1). Note that the extension $R\to R^{\prime}$ does not change when we shrink $\tau^{\prime}$, as it follows from the construction.
@@
For a higher $\Theta$-stratum of the cone $\tau$ on an algebraic stack $\mathcal{M}$, which we denote as $\mathcal{Z}^{+}\subset{\rm Map}(\Theta_{\tau},\mathcal{M})$, consider restriction map $({\rm exp}\otimes\iota)^{*}\colon\mathcal{Z}\to{\rm Map}(\Theta_{\tau},\mathcal{M})$ and suppose it gives the $\Theta$-strata structure for each $\xi$ i.e., the images are connected. 444at the moment, we are not sure how weak or strong this condition is Then we call the obtained $\Theta$-strata as restriction along $\xi$ of the $\Theta$-strata $\mathcal{Z}$ and often denote as $\mathcal{Z}_{\xi}$.
@@ 3.3. Stable reduction along higher $\Theta$-stratification
    a totally ordered 555our order is converted from the notation in [HL14] for convenience to study problems related to K-stability, as in [BHLLX21] set $\Gamma$, which has a maximal element $1$,
@@
    closed666it is a priori not necessarily reduced, although in our stable reduction type theorem, the scheme structure is not really important substacks $\mathcal{M}_{<c}$ of $\mathcal{M}$ for $c\in\Gamma$ which satisfies that
@@
  • in particular, $f_{l}((\mathfrak{m}_{l})\in\mathcal{M}_{=1}$.
@@ Acknowledgement
The author thanks R.Conlon, E.Inoue, M.Jonsson for the comments on the manuscript. During the research, the author was partially supported by Grant-in-Aid for Scientific Research (B) 23H01069, 21H00973, Grant-in-Aid for Scientific Research (A) 21H04429, 20H00112, and Fund for the Promotion of Joint International Research (Fostering Joint International Research) 23KK0249. Part of this work is done during the author’s stay at Institut de mathématiques de Jussieu - Paris Rive Gauche. We appreciate the kind hospitality.

### v2

@@ 3.2. Higher $\Theta$-stable reduction theorem
We take a simplicial subcone $\sum_{i=1}^{r}\mathbb{R}_{\geq 0}v^{\prime}_{i}$ of $\tau^{\prime}$, where $\{v^{\prime}_{i}\}_{i}\cup(\vec{0},1)$ is a basis of of $N$, which still contains $\xi^{\prime}$. We also take the cone which contains $(\xi,[c^{\prime}-\epsilon,c^{\prime}])$ for $0<\epsilon\ll 1$ and denote it as $\widetilde{\tau}^{\prime}$. Write $v^{\prime}_{i}=(v_{i},\frac{l_{i}}{m_{i}})$ with $l_{i},m_{i}\in\mathbb{Z}_{>0}$ and set $m:={\rm lcm}\{m_{i}\}$.
Take an extension of $R$ as follows. If $R$ is a discrete valuation ring, then consider $R\to R^{\prime}$ with ramification degree $m$, which is obtained as a localization of $R[x]/(x^{m}-\pi)$ at $(x)$. Note that $\{\frac{m_{i}v^{\prime}_{i}}{m}\}_{i}$s together with $(\vec{0},\frac{1}{m})$ gives a basis of a regular cone of $N\oplus\frac{1}{m}\mathbb{Z}$. From the construction, (base change of) $\tilde{f}$ is extendable to $\widetilde{\rm ST}_{R^{\prime}}(\tau^{\prime})\to M$ and ${\overline{\rm ST}}_{R^{\prime}}(\tau^{\prime})\to\mathcal{M}$ and the desired properties are satisfied from the construction and finish the proof of (1). Note that the extension $R\to R^{\prime}$ does not change when we shrink $\tau^{\prime}$, as it follows from the construction. If $R$ is not divisorial valution ring, then we instead apply Lemma 3.5.
@@
For a higher $\Theta$-stratum of the cone $\tau$ on an algebraic stack $\mathcal{M}$, which we denote as $\mathcal{Z}^{+}\subset{\rm Map}(\Theta_{\tau},\mathcal{M})$, consider restriction map $({\rm exp}\otimes\iota)^{*}\colon\mathcal{Z}\to{\rm Map}(\Theta_{\tau},\mathcal{M})$ and suppose it gives the $\Theta$-strata structure for each $\xi$ i.e., the images are connected. 555at the moment, we are not sure how weak or strong this condition is Then we call the obtained $\Theta$-strata as restriction along $\xi$ of the $\Theta$-strata $\mathcal{Z}$ and often denote as $\mathcal{Z}_{\xi}$.
@@ 3.3. Stable reduction along higher $\Theta$-stratification
    a totally ordered 666our order is converted from the notation in [HL14] for convenience to study problems related to K-stability, as in [BHLLX21] set $\Gamma$, which has a maximal element $1$,
@@
    closed777it is a priori not necessarily reduced, although in our stable reduction type theorem, the scheme structure is not really important substacks $\mathcal{M}_{<c}$ of $\mathcal{M}$ for $c\in\Gamma$ which satisfies that
@@
  • in particular, $f_{l}(\mathfrak{m}_{l})\in\mathcal{M}_{=1}$.
@@ Acknowledgement
The author thanks E.Inoue, M.Jonsson for the comments on the manuscript. During the research, the author was partially supported by Grant-in-Aid for Scientific Research (B) 23H01069, 21H00973, Grant-in-Aid for Scientific Research (A) 21H04429, 20H00112, and Fund for the Promotion of Joint International Research (Fostering Joint International Research) 23KK0249. Part of this work is done during the author’s stay at Institut de mathématiques de Jussieu - Paris Rive Gauche. We appreciate the kind hospitality.